• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name eschulte

    Log entry time 08:12:19 on February 10, 2007

    Entry number 191993

    keyword=Saturday 2/10 Owl Shift Log

    All Run Numbers are referenced from HRS-L

    00:00 Shift starts with "Y" component of raster not working properly. 00:46 Back to debugging the raster problem after some fussing over the "lithinum wierdness". Request 5 uA, 3x3 (14x4 from MCC) raster size on C target. Run# 6109. Raster for 6109 looks better now! It's a miracle! 6110 also seems to be OK.

    Please note: The pre-scale factors for runs 6109 through 6111 were not set to minimuze the dead-time (it was around 60%).

    01:40 Changed target to "5" (supposed lithium) to check the "lithium wierdness" spectrum.

    01:45 At 2uA we noticed same raster problem as earlier in shift.

    01:50 Resume running on carbon target with 2 uA to check raster problem.

    01:54 Change to empty target to test our target ladder position effect on raster behavior theory (i.e. the farther "up" the ladder is the more messed up the raster).

    02:01 Resume running with the empty target to check the raster & other odd behavior related to the lithium target. We're reasonably convinced that lithium is in fact in the lithium target now.

    02:10 Raster looks fine with target empty, so what's up with lithium?

    02:15 Request change to lithium, we'll try with the pulser on to se if we a "fantom" raster problem. Encoder reported being "off" by 3 "encoder units" but in the correct readback position for the carbon target. Back to lithium. Now it's off by 14 "encoder units".

    02:31 Despite the pulser included the raster still looks odd on the lithium target; however, with the pulser only it looks OK.

    02:41 Request beam spot move +0.5mm BPM-B Y

    02:56 Request beam spot move of another +0.5 mm on BPM-B Y to a total of +1mm

    03:22 Request beam spot move to a total of -1mm on BPM-B Y.

    03:26 Rates consistant with the empty target position. The on-line profiles and spot++ suggest that we don't know where the lithium target is. We'll go back to carbon, tantalum, & boron carbide to 2M events each. Recentering beam on target.

    03:42 Upon re-centering carbon we noticed that, under the assumption that all target centers should be in the same location on the monitor, that the lituium target moves to a position an entire target-size off. Bummer.

    03:50 Resume taking data with carbon target. Reset pre-scales to a dead time around 10% (1-2% for 6122).

    04:43 Request change to tantalum target. Target looks the same when in position as the carbon target. Momentum (delta) looks about the same with the 3x3 raster as with the smaller one used earlier.

    05:30 Continue with Ta production data. Beem has been excellent.

    06:12 Request target change to boron carbide. MCC reports that they will be doing a "QE" during the target change.

    06:17 Boron carbide target in position & ready to take beam back when MCC is ready to deliver it. When in position, the Boron carbide target matches well the C & Ta target positions.

    06:40 Gilman requests the current be dropped to 1uA to reduce the real trigger rate. We'll see if the fast feedback will keep working at this low current.

    06:49 Restart taking data with 1uA beam. Run #6128. Beam position a little more "ragged" at this lower current, but well within limits.

    07:15 Paged Brad (RC) about impending target change. Will probably do Gilman's current scan to fill up the time until we get the Li target encoder position fixed. When Brad called back, we decided on the scan and that he would give Dave Meekins a call to get the encoder problem corrected.

    07:21 Changed current to 2uA for current scan. Will take 1 4M event file. Run #6130 Using prescale factors for Run #6129.

    07:44 Changed current to 3uA for continuation of current scan. Will take 1 4M event file. Run #6131. Using same prescale factors as for Run #6127.

    08:00 Shift ended with a report from Brad (RC) with info about what to change the encoder value to; however, we don't know how to change this in the GUI. Day shift paged D. Meekins via Brad for more information on how to update the GUI.